Question 41
ProbabilityA box contains black, white and red identical balls. The probability of picking a black ball at random from the box is 3/10 and the probability of picking a white ball at random is 2/5. If there are 30 balls in the box, how many of them are red?
- A.3
- B.7
- C.9
- D.12
Show the answer and explanation
Answer: C (9)
Explanation
P(red) = 1 − 3/10 − 4/10 = 3/10, so there are 3/10 × 30 = 9 red balls.